suppose the united states is currently producing 200 tons of hamburgers and 90 tons of tacos mexico is cutrently producing 20 tons of hamburgers and 50 tons of tacoes if the united states and mexico each specialize in producing only one good the good for which each has a comparitive advantage then a toal of ? additional tons of hamburgers can be produced for the two countries combines and a total of ? additional tons of tacoes can be produced
sented in the table below. United States Mexico Hamburgers (in tons) Tacos (in tons) Hamburgers (in tons) Tacos (in tons) 270 150 100 180 10 100 200 90 20 50 300 0 30 0 tons of tacos. If the United States and Mexico each specialize in producing only one good (the good for which each has a c
This Question: 1 pt Suppose the United States and Mexico both produce hamburgers and tacos. The combinations of the two goods that each country can produce in one day are presented in the table below. United Stater Hamburgers (in tons) C 100 200 300 Which country has an absolute advantage in producing tacos? The United States Which country has a comparative advantage in producing tacos? Mexico Suppose the United States is curently producing 200 tons of hamburgers and 90 tons of tacos and Mexico is currenty producing 20 tons of hamburgers and 50 tons of tacos. If the United States and an integer) and a total of additional ton(s) of tacos can be produced.
Answer #1

Q) Which country has an absolute advantage in producing tacos = The United States

When the US produces 0 Hamburgers, it can produce 270 tacos and when Mexico produces 0 Hamburgers, it can produce 150 tacos. Since, US can produce more tacos, it has a comparative advantage in tacos.

Q) Which country has a comparative advantage in producing tacos = Mexico

Lets see who has a comparative advantage in producing which good:

Hamburgers Tacos
US 100 180
Mexico 10 100

We see from the table below that 1 Hamburger costs = 1.8 Tacos for the US, 1 Taco costs = 0.55 Hamburger for the US. Similarly, 1 Hamburger costs = 10 Tacos for Mexico, 1 Taco costs = 0.1 Hamburger for Mexico. SInce. 1.8<10, US has a comparative advantage in producing Hamburgers and since 0.1<0.55, Mexico has a comparative advantage in producing Tacos.

Hamburgers Tacos
US 100 = 180/100 = 1.8 180 = 100/180 = 0.55
Mexico 10 = 100/10 = 10 100 = 10/100 = 0.1

Q) Suppose the United States is currently producing 200 tons of hamburgers and 90 tons of tacos and Mexico is currently producing 20 tons of hamburgers and 50 tons of tacos. If the United States and Mexico each specialize in producing only one good for which each has a comparative advantage then a total of 80 additional tons of hamburgers can be produced when the two countries combines and a total of 10 additional tons of tacos can be produced.

If the United States specializes only in producing Hamburgers, it can produce a total of 300 hamburgers as given in your graph. Similarly, if Mexico specializes only in producing Tacos, it can produce a total of 150 tacos as given in your graph.

So, if the United States is currently producing 200 tons of hamburgers and 90 tons of tacos and Mexico is currently producing 20 tons of hamburgers and 50 tons of tacos, a total of 200 + 20 = 220 hamburgers and 90 + 50 = 140 tacos is being produced.

f the United States and Mexico each specialize in producing only one good for which each has a comparative advantage then a total of 300 - 220 = 80 additional tons of hamburgers can be produced when the two countries combines and a total of 150 - 140 = 10 additional tons of tacos can be produced.